OpenStudy (anonymous):

A rocket burns up at a rate of 3 tons/min after falling out of orbit into the atmosphere. if the rocket weighed 5100 tons before rentry, express its weight w as a function of the time T, in mins, or rentry.

OpenStudy (anonymous):

im not doing so well with these funtion problems

OpenStudy (anonymous):

i try

OpenStudy (anonymous):

w=5100-3T rearrange T=(5100-w)3

OpenStudy (anonymous):

tell me if this wrong :)

OpenStudy (anonymous):

nice checks out for me thank you

OpenStudy (anonymous):

wait2 sorry it should be T=(5100-w)/3